Chuck exerts a force on a 1.1 kg push broom as he walks across the floor at a constant speed. The coefficient of friction between the floor and the broom is 0.45 and the broom handle makes an angle of 41 with the horizontal. Determine the amount of force with which Chuck pushes downward (along the handle of the broom) to achieve this constant speed motion


1
Expert's answer
2019-09-30T10:21:19-0400
"N=mg+T\sin{41\degree}"

"\mu N=T\cos{41\degree}"

"\mu (mg+T\sin{41\degree})=T\cos{41\degree}"

"0.45 ((1.1)(9.8)+T\sin{41\degree})=T\cos{41\degree}"

"T=11\ N"